OpenStudy (anonymous):

what are the zeros of h(x)=2x^2-16x+27

OpenStudy (anonymous):

As it's a quadratic equation you can simply use this formula: \[ x_{1,2} = \frac{1}{2a}\left[-b\pm \sqrt{b^2-4ac}\right] \]

OpenStudy (anonymous):

Do you know it?

OpenStudy (anonymous):

to find the possible zeros?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

it's basically the standard formula

OpenStudy (anonymous):

but i think there is more then 1 zero

OpenStudy (anonymous):

okay let me show you

OpenStudy (anonymous):

thank you because i dont understand

OpenStudy (anonymous):

??

OpenStudy (anonymous):

\[h(x) = \underbrace{2}_{a}x^2-\underbrace{16}_{b}x+\underbrace{27}_c\] Now we plug that into the formula:\[ x_{1/2} = \frac{1}{4}\left[ -(-16) \pm \sqrt{16^2 - 4\cdot2\cdot27}\right] \] \[ = \frac{1}{4}\left[ 16 \pm \sqrt{256 - 216}\right] \] \[ = \frac{1}{4}\left[ 16 \pm \sqrt{40}\right] \] \[ = \frac{1}{2}\left[ 8 \pm 2\sqrt{10}\right] \] Now we can get the zeros as:\[ x_1 = \frac{1}{2}\left[ 8 + 2\sqrt{10}\right], \quad x_2 = \frac{1}{2}\left[ 8 - 2\sqrt{10}\right] \]

OpenStudy (anonymous):

i got the same answer, thanks so much

OpenStudy (anonymous):

You're right that there are two solutions, they exists because we take the squareroot of something. Which means there are necessarily two solutions, because the result of that squareroot can either be negative or positive. \[ \sqrt{-a^2} = \sqrt{-a(-a)} = \sqrt{a^2} = a \] So we cannot tell whether which means we cannot tell whether a was positive or negative int the beginning.

OpenStudy (anonymous):

You're welcome

OpenStudy (anonymous):

how do i find the y- intercept of f(x)= (x^2-6x-7)/(4x^2-36x+56)